OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, –5) with a slope of –1?

OpenStudy (anonymous):

y = mx + b... plug in the point given to get the y-interept... -5 = (-1)(1) + b.. -4 = b.. so the equation would be y=x-4